OpenStudy (anonymous):

please help medal and fan For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8

OpenStudy (anonymous):

-2y + 6 2y - 6 2y + 6 -2y - 6

OpenStudy (anonymous):

do u jst need to solve using subsitution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with these 2y - 6 2y + 6 -2y - 6

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yes...confused though

OpenStudy (anonymous):

so am I do u know any one that can help

OpenStudy (anonymous):

@TheSmartOne PLEASE help

OpenStudy (anonymous):

do u understand this

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

@mathmate please help

TheSmartOne (thesmartone):

I got to leave sorry. But @mathmate @sleepyjess @pooja195 might be able to help you.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do u understand this

OpenStudy (mathmate):

Given: -x + 2y = -6 3x + y = 8 From first equation: -x + 2y = -6 2y+6=x or \(\color{red}{x}=\color{red}{(2y+6)}\) Substitute into second equation: \(3(\color{red}{2y+6} )+ y = 8\) Distribute the multiplication by 3 and solve for y.

OpenStudy (anonymous):

so it would be 6y+18+y=8?

OpenStudy (mathmate):

Yes, so far so good!

OpenStudy (anonymous):

ok now what do I do

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Isolate y on the left, and numbers on the right, and solve for y. Y will be a fraction, but just keep going.

OpenStudy (anonymous):

so it would u combine like terms to

OpenStudy (mathmate):

To isolate y, you need to first add together the like terms of y: 6y+18+y=8 becomes 7y+18=8 Can you continue?

OpenStudy (anonymous):

yes ok so it would then be 18=8-7y

OpenStudy (mathmate):

If I were you, I would keep the 7y on the left, but subtract 18 from both sides. You want to keep variables on the left, and numbers on the right. 7y+18=8 subtract 18 from both sides 7y+18-18=8-18 Can you continue?

OpenStudy (anonymous):

ok so it would be 7y=-10 and no

OpenStudy (mathmate):

Don't worry, as I said, the answer for y will not be a whole number. I have to go now, perhaps you can tag others to continue. You're doing fine so far.

OpenStudy (anonymous):

wait can I just tell you what I think the answer is

TheSmartOne (thesmartone):

I am here now xD

OpenStudy (anonymous):

yay thank you

OpenStudy (anonymous):

if you look at the choices I dont think were doing this right -2y + 6 2y - 6 2y + 6 -2y - 6

TheSmartOne (thesmartone):

Hmm well the question didn't ask to solve for a variable..

OpenStudy (anonymous):

right

TheSmartOne (thesmartone):

It says to isolate x in -x + 2y = -6

OpenStudy (anonymous):

yeah so do I just subtract 2y from both sides

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @mathmate Given: -x + 2y = -6 3x + y = 8 From first equation: -x + 2y = -6 2y+6=x or \(\color{red}{x}=\color{red}{(2y+6)}\) Substitute into second equation: \(3(\color{red}{2y+6} )+ y = 8\) Distribute the multiplication by 3 and solve for y. \(\color{blue}{\text{End of Quote}}\) The answer is in red.. lol

TheSmartOne (thesmartone):

But we have -x + 2y = -6

OpenStudy (anonymous):

Oh I feel stupid thank you

TheSmartOne (thesmartone):

We could just add x and add 6 on both sides.

OpenStudy (anonymous):

so thats not the answer

TheSmartOne (thesmartone):

-x + 2y = -6 +x +x ------------- 2y= -6 + x 2y = -6 + x +6 +6 ----------- 2y+6 = x

TheSmartOne (thesmartone):

It is.

OpenStudy (anonymous):

of thank you

OpenStudy (anonymous):

oh*

TheSmartOne (thesmartone):

np :)

OpenStudy (anonymous):

can u help me with another one please